# CoCoST: Automatic Complex Code Generation with Online Searching and Correctness Testing

### v2

@@ 1 Introduction
Figure 1: An Example of the Human Developer Code-writing Process Imitated by the CoCoST. After the problem is received, an online search is performed to simulate search results and create an initial version of code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
@@
To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be publicly available upon acceptance. (Automatic Complex Code Generation with Online Searching and Correctness Testing) that improves the generation and refinement of complex code by LLMs through the planned online searching and automatic correctness testing steps. The intuition of CoCoST is straightforward: During the coding process, most human developers are not bothered by the above challenges, as illustrated in Figure 1. Developers can easily overcome these obstacles by searching online through engines (e.g., Google and Bing) for solutions, experiences, and guidelines. In addition, they can create test cases and execute code to ensure the correctness of the code logic.
@@
To address Challenges 1, CoCoST proposes an online search methodology. This process involves querying web search engines and then extracting pertinent information to construct prompts for LLMs. The approach presents several benefits: (1) Retrieving information from the up-to-date blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. (2) Online search extends beyond the scope of static offline documentation, covering a wider array of problems without being confined to a predetermined set. Meanwhile, it reduces the effort developers need to expend in assembling documentation, thereby increasing the framework’s level of automation. However, using problem descriptions as search queries can be challenging due to the complexity of the problems, which often encompass multiple issues. Therefore, we propose an online search with generation query through planning.
@@
To address the Challenge 2, we introduce generation of test cases during refinement. Several studies Chen et al. (2022); Shinn et al. (2023) have attempted to generate tests. However, these methods often fall short when applied to the generation of complex code due to its intricate logic and outputs, which complicate the direct production of accurate tests (both the inputs and expected outputs for the solution code). CoCoST utilizes LLMs to automatically generate test cases (the inputs for the code). This strategy cleverly focuses on generating test cases without attempting to produce complete tests. It significantly simplifies the process of test case generation and facilitates its precise creation for complex code.
@@
To address the Challenge 3, this work prioritizes correctness testing in refinement. During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors. CoCoST incorporate both the execution output results and the errors into within the refinement prompts for LLMs to enhance the correctness. Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, Matplotlib charts. Thus, CoCoST propose serialization of input and output to convert them into understandable sequences before being processed by LLMs. Particularly those are excessively long or non-textual modalities.
@@
We evaluate the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval). By comparing with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and average 9.47% on ClassEval.Moreover, we analyze and discovery that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem.
In summary, our main contributions are:
@@
  • We propose the novel CoCoST framework to generate complex code, and can be automatic in real-world scenarios.
@@
  • We apply online search in code generation for the first time to our knowledge, in order to generate the complex code.
@@
  • We prioritize correctness testing in refinement with test case generation and serialization of input and output, in order to refine hidden bugs in complex code.
@@
  • We conduct experiments on DS-1000 and ClassEval datasets anddemonstrate the effectiveness and universality of CoCoST.
@@ 2 Related Work
Retrieval-augmented code generation. With the emergence of Large Language Models (LLMs), a variety of retrieval-augmented techniques have been developed to compensate for issues such as the inherent knowledge limitations. DocPromptZhou et al. (2023) and SELVEVOLVEJiang et al. (2023) leverage document libraries or models as knowledge bases to improve code generation. However, their reliance on fixed document libraries limits the scope of information they can provide and confines the generated code to the context of these libraries. Furthermore, the prerequisite of preestablished document libraries prevents these approaches from being fully autonomous in real-world frameworks. Solutions such as WebGPT Nakano et al. (2022), LaMDA Thoppilan et al. (2022), and FreshLLMs Vu et al. (2023) enhance the performance of natural language tasks by using online search or open web knowledge. However, because complex code generation often involves multiple steps and complexities, these methods struggle with direct application to complex code generation.
Code refinement. Refine iteratively enhances generated code for greater precision. Self-DebugChen et al. (2023), SELFEVOLVEJiang et al. (2023), and Self-EditZhang et al. (2023) improve code generation by refining code through the resolution of errors identified during execution. These methods effectively address errors, while when it comes to complex code generation, subtle bugs also play a significant role in the overall error landscape. Moreover, relying on pre-existing tests from datasets in refinement limits their autonomy in real-world applications, where such tests may not be readily available. CodeT Chen et al. (2022), Reflexion Shinn et al. (2023), and CODECHAIN Le et al. (2023) seek to strengthen code generation by creating tests. But the tests they generate include not only the inputs for the solution code but also the expected outputs. This poses a substantial challenge for complex code generation, where the logic can be intricate and certain problems may not lend to straightforward ground truth generation.
Figure 2: The Pipline of CoCoST. Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps. These queries enable the retrieval of diverse information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases for testing the initial code. The test results serve as crucial inputs for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
@@ 3 Methodology
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieval, where relevant information is obtained through online search and utilized by LLMs to generate initial code. The second step is refinement, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
@@ 3.1.1 Generation Query through Planning
In order to generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem. The planning phase involves outlining the natural language steps $P=\{plan_{1},\ldots,plan_{N}\}$ required to address the problem. Later, the assessment involves utilizing LLMs to determine whether each planning step requires an online search. Subsequently, the planning steps identified as necessitating online search are translated into queries $Q=\{q_{1},\ldots,q_{N}\}$ for use in the subsequent search process.
@@ 3.1.2 Online Search
For the above generated queries, we conduct an online search. In this study, we use the online search API [^2] https://github.com/Nv7-GitHub/googlesearch for the search process as Equation (3). CoCoST can also be applied to private or domain-specific knowledge repositories as long as they are accessible via query, with details in §A.
@@ 3.2.2 Generation of Test Cases
CoCoST introduces generation of test cases with LLMs to adapt to real-world scenarios.
@@ 4.1.1 Datasets
We conduct experiences on two complex code generation datasets:
@@ 4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CoCoST surpasses the current SOTA framework, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CoCoST exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CoCoST in generating complex code. CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refining the details to further enhance the correctness of the code.
@@ 4.3 Ablation Study
CoCoST significantly enhances the base model’s ability to generate complex code. Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
@@
Online search outperforms offline retrieval in effectiveness and has a wider range of applicability. As shown in Table 1, using only online retrieval (the row w/o generation of test case) outperforms DocPrompting, which is an offline retrieval approach. Moreover, in real-world scenarios, as opposed to specific datasets, the types of problems encountered are more diverse. The scalability of online retrieval enables them to effectively address a wide range of problems. However, offline retrieval systems struggle to encompass all relevant information comprehensively.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval dataset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
@@ 4.4 Analysis of Different Base Models Performance
Code generation necessitates models to have in-context learning abilities. The generated code should be built on all the above provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder is comparatively interior to GPT models. This could be due to WizardCoder’s limited in-context learning ability, especially with the complex and lengthy prompts, hindering accurate context comprehension and code modification.
4.5 Case Study and Error Analysis
For the case study on online retrieval, refer to Figure 4. It can be observed that by imitating the usage of functions found through online search, the model is better equipped to prepare the required parameters for the functions and to generate corresponding code. This significantly reduces the difficulty of generating complex code.
For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thus improving hidden errors and generating the correct code.
Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors. For the former, a discussion is provided in §C.2, while for the latter, an error analysis is conducted in §C.3.
@@ 5 Conclusion
In conclusion, CoCoST introduces a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CoCoST outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
@@
The main limitation of our research is that it has underlying issues of exceeding the allowed times of accesses due to multiple calls to the Google Search API. Similarly, we also have made multiple API calls to test and enhance the performance of the GPT models.
Ethics Policy
This research does not pose any ethical concerns. The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
@@ C.1 DS-1000 Results
    C.2 Analysis of Pipeline
    Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval. On the DS-1000 dataset using GPT-4, this approach result in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%). This confirms that the generated plans are significantly effective.
    C.3 Error Analysis
    For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements. Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect. For the former cases, it shows that CoCoST is able to fix some bad cases in the refinement stage even though the retrieval contents have some errors. For the later cases, we do observe some cases that are worthy of further research as potential directions for future work. First, the search content could be more detailed. E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but is exactly where the bug in the code. Second, the search query could be more targeted. The descriptions of some queries are not specific enough in terms of some complicated problems. Therefore, more sub-queries are needed to help model receive clearer instructions.
C.4 Case Study
Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset. Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices. After unsuccessful attempts, the model turns to an online search through a self-generated query and integrates the research results as part of the next prompt. Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class. Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
Figure 3: Case Study for Correctness Testing.
Figure 4: Case Study for Online Retrieval.

### v3

@@ 1 Introduction
Figure 1: An Example of the Human Developer Code-writing Process Imitated by the CoCoST. After the problem is received, an online search is performed to simulate search results and create an initial version of the code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
@@
To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be open-sourced on https://github.com/microsoft/CoCoST. (Automatic Complex Code Generation with Online Searching and Correctness Testing) that improves the generation and refinement of complex code by LLMs through the planned online searching and automatic correctness testing steps. The intuition of CoCoST is straightforward: During the coding process, most human developers are not bothered by the above challenges, as illustrated in Figure 1. Developers can easily overcome these obstacles by searching online through engines (e.g., Google and Bing) for solutions, experiences, and guidelines. In addition, they can create test cases and execute code to ensure the correctness of the code logic.
@@
To address Challenge 1, CoCoST proposes an online search methodology. This process involves querying web search engines and then extracting pertinent information to construct LLM prompts. The approach presents several benefits: (1) Retrieving information from the up-to-date blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. (2) Online search extends beyond the scope of static offline documentation, covering a wider range of problems without being confined to a predetermined set. Meanwhile, it reduces the effort developers need to expend in assembling documentation, thereby increasing the framework’s level of automation. Using problem descriptions as search queries can be difficult, because problems are generally intricate and include several components. Therefore, we propose an online search with query generation through planning.
@@
To address Challenge 2, we introduce generation of test cases during refinement. Several studies Chen et al. (2022); Shinn et al. (2023) have attempted to generate tests. However, these methods often fall short when applied to the generation of complex code due to its intricate logic and outputs, which complicate the direct production of accurate tests (both the inputs and expected outputs for the solution code). CoCoST utilizes LLMs to automatically generate test cases (the inputs for the code). This strategy cleverly focuses on generating test cases without attempting to produce complete tests. It significantly simplifies the process of test case generation and facilitates its precise creation for complex code.
@@
To address Challenge 3, this work prioritizes correctness testing in refinement. During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking the existence of the errors. CoCoST incorporates both the execution output results and the errors within the refinement prompts for LLMs to enhance the correctness. Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, and Matplotlib charts. Thus, CoCoST proposes serialization of input and output to convert them into understandable sequences before being processed by LLMs. Particularly those are excessively long or non-textual modalities.
@@
We evaluated the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval). Compared with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and an average of 9.47% on ClassEval.Moreover, we analyze and discover that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem. In summary, our main contributions are as follows.
@@
  • We propose the novel CoCoST framework to generate complex code. CoCoST can be automatic in real-world scenarios.
@@
  • To generate complex code, we designed an online search method (query generation) in code generation for the first time to our knowledge.
@@
  • To refine hidden bugs in complex code, we prioritize correctness testing in refinement with test case generation and serialization of input and output data types.
@@
  • We conducted experiments on the DS-1000 and ClassEval datasets to demonstrate the effectiveness and universality of CoCoST.
@@ 2 Related Work
Retrieval-augmented code generation.
With the emergence of Large Language Models (LLMs), a variety of retrieval-augmented techniques have been developed to compensate for issues such as the inherent knowledge limitations. DocPrompt Zhou et al. (2023) and SELVEVOLVE Jiang et al. (2023) leverage document libraries or models as knowledge bases to improve code generation. However, their reliance on fixed document libraries limits the scope of information they can provide and confines the generated code to the context of these libraries. Furthermore, the prerequisite of preestablished document libraries prevents these approaches from being fully autonomous in real-world frameworks. Solutions such as WebGPT Nakano et al. (2022), LaMDA Thoppilan et al. (2022), and FreshLLMs Vu et al. (2023) enhance the performance of natural language tasks by using online search or open web knowledge. However, because complex code generation often involves multiple steps and complexities, these methods struggle with direct application to complex code generation.
Code refinement.
Refine iteratively enhances generated code for greater precision. Self-Debug Chen et al. (2023), SELFEVOLVE Jiang et al. (2023), and Self-Edit Zhang et al. (2023) improve code generation by refining code through the resolution of errors identified during execution. These methods effectively address errors, while when it comes to complex code generation, subtle bugs also play a significant role in the overall error landscape. Moreover, relying on pre-existing tests from datasets in refinement limits their autonomy in real-world applications, where such tests may not be readily available. CodeT Chen et al. (2022), Reflexion Shinn et al. (2023), and CODECHAIN Le et al. (2023) seek to strengthen code generation by creating tests. But the tests they generate include not only the inputs for the solution code but also the expected outputs. This poses a substantial challenge for complex code generation, where the logic can be intricate and certain problems may not lend to straightforward ground truth generation.
Figure 2: The Pipline of CoCoST. Step 1: LLM is employed to strategize the problem and formulate queries based on the outlined steps. These queries enable the retrieval of various information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases to test the correctness of the initial code. The serialization of test results serves as crucial input for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
@@ 3 Methodology
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieval, where relevant information is obtained through an online search and utilized by LLMs to generate initial code. The second step is refinement, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
@@ 3.1.1 Generation Query through Planning
To generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem. The planning phase involves outlining the natural language steps $P=\{plan_{1},\ldots,plan_{N}\}$ required to address the problem. Later, the assessment involves utilizing LLMs to determine whether each planning step requires an online search. Subsequently, the planning steps identified as necessitating online search are translated into queries $Q=\{q_{1},\ldots,q_{N}\}$ for use in the subsequent search process.
@@ 3.1.2 Online Search
For the above-generated queries, we conduct an online search. In this study, we use the online search API [^2] https://github.com/Nv7-GitHub/googlesearch for the search process as Equation (3). CoCoST can also be applied to private or domain-specific knowledge repositories as long as they are accessible via query, with details in §A.
@@ 3.2.2 Generation of Test Cases
CoCoST introduces a generation of test cases with LLMs to adapt to real-world scenarios.
@@ 4.1.1 Datasets
We conduct experiences on two complex code-generation datasets:
@@ 4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CoCoST surpasses the current SOTA framework, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CoCoST exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CoCoST in generating complex code. CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code and refining the details to further enhance the correctness of the code.
@@ 4.3 Ablation Study
CoCoST significantly enhances the base model’s ability to generate complex code. Compared to the base model, CoCoST has shown improvements of 7.8% on the DS-1000 dataset and an average of 9.47% on ClassEval, demonstrating the effectiveness of the CoCoST.
@@
Online search outperforms offline retrieval in effectiveness and has a wider range of applicability. As shown in Table 1, using only online retrieval (the row w/o generation of the test case) outperforms DocPrompting, which is an offline retrieval approach. Moreover, in real-world scenarios, as opposed to specific datasets, the types of problems encountered are more diverse. The scalability of online retrieval enables them to effectively address a wide range of problems. However, offline retrieval systems struggle to encompass all relevant information comprehensively.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CoCoST shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and an average of 2.7% and 0.3% on the ClassEval dataset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
@@ 4.4 Analysis of Different Base Models Performance
Code generation necessitates models to have in-context learning abilities. The generated code should be built on all the above-provided contents, and the understanding of the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder is comparatively interior to GPT models. This could be due to WizardCoder’s limited in-context learning ability, especially with the complex and lengthy prompts, hindering accurate context comprehension and code modification.
4.5 Cascade Analysis
Our framework consists of multiple components cascaded together, which results in certain intermediate steps that cannot be explicitly validated for effectiveness, as well as the potential generation of cascading errors. For the former, a discussion is provided in §4.5.1, while for the latter, an error analysis is conducted in §4.5.2.
4.5.1 Analysis of Pipeline
Regarding generating plans, we attempt to bypass the planning step and directly perform online retrieval. On the DS-1000 dataset using GPT-4, this approach results in a performance of 55.70%, which is 6.4% lower than using planning for online retrieval (62.10%), and even 4.5% lower than only using GPT-4 to generate code (60.20%). This confirms that the generated plans are significantly effective.
Figure 3: Case Study for Correctness Testing.
4.5.2 Error Analysis
For cascading errors, some errors generated by retrieval processes can be corrected through refinement, while others may persist, necessitating future improvements. Take DS-1000 as an example: Compared to the baseline (GPT-4), we observed that among the instances that turned erroneous after retrieval, 39.6% were corrected, while 60.4% remained incorrect. For the former cases, it shows that CoCoST can fix some bad cases in the refinement stage even though the retrieval contents have some errors. For the later cases, we do observe some cases that are worthy of further research as potential directions for future work. First, the search content could be more detailed. E.g., Some basic steps that LLMs consider unnecessary to search for are not generating queries, but are exactly where the bug is in the code. Second, the search query could be more targeted. The descriptions of some queries are not specific enough in terms of some complicated problems. Therefore, more sub-queries are needed to help the model receive clearer instructions.
Figure 4: Case Study for Online Retrieval.
4.6 Case Study
For the case study on online retrieval, refer to Figure 4. It can be observed that by imitating the usage of functions found through online search, the model is better equipped to prepare the required parameters for the functions and to generate corresponding code. This significantly reduces the difficulty of generating complex code. For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thus improving hidden errors and generating the correct code.
@@ 5 Conclusion
In this paper, we propose CoCoST, a novel framework for generating complex code in real-world scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CoCoST outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
@@
The primary limitation of our research is exceeding the allowed access limits due to frequent calls to the Google Search API. Likewise, we made multiple API calls to test and optimize the performance of the GPT model.
Ethics Statement
The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work. All the datasets used in this paper have been reviewed to ensure that they do not contain personally identifiable information or offensive content. However, since these datasets are sourced from the Internet, potential bias may still be present. Furthermore, despite our careful review, the process involving the LLMs may inadvertently introduce inappropriate information into the evolved data.
Acknowledgments
We thank all anonymous reviewers for their valuable comments. Xinyi He and Zejian Yuan were supported in part by the National Key R&D Program of China (2023YFB4704900) and NSFC (61976170, 62088102).